OpenStudy (anonymous):

Jose invests $3, 250 at 6% interest compounded annually. What will be the balance in the account after 3.5 years?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

ohhhhh thank you!

OpenStudy (campbell_st):

well its simply a case of \[A = 3250(1 + 0.06)^{3.5}\] where A is the future value of the investment

OpenStudy (anonymous):

Ok I will calculate it.

OpenStudy (anonymous):

ugg I got 12057.5

OpenStudy (anonymous):

What did I do wrong? @campbell_st

OpenStudy (whpalmer4):

okay, what do you get for 1+0.06? then what do you get for that quantity to the 3.5 power?

OpenStudy (anonymous):

Yeah! Whpalmer4!

OpenStudy (anonymous):

1.06

OpenStudy (whpalmer4):

good, that's what I expected :-)

OpenStudy (anonymous):

I even tried 3250 (1.06)^3.5

OpenStudy (whpalmer4):

well, \[3250*(1.06)^{3.5}\]IS what you want to calculate

OpenStudy (whpalmer4):

why don't you try just calculating \(1.06^{3.5}\)

OpenStudy (anonymous):

I did this on my calculator : 1.06 *3.05 = X then X * 3250

OpenStudy (anonymous):

ok I will calculate what you typed. Just sec

OpenStudy (whpalmer4):

No, you have to raise 1.06 to the 3.5 power.

OpenStudy (anonymous):

3.71 so far

OpenStudy (whpalmer4):

it's the same as multiplying \(1.06*1.06*1.06*\sqrt{1.06}\)

OpenStudy (whpalmer4):

What are you using for a calculator?

OpenStudy (whpalmer4):

You're just multiplying, not doing exponentiation

OpenStudy (anonymous):

basic calculator. I got 3.18 now

OpenStudy (whpalmer4):

Okay, that won't work. You need something that has exponentiation. You can use google: https://www.google.com/#q=1.06%5E3.5&safe=off

OpenStudy (whpalmer4):

you type in 1.06, then press the \(y^x\) button, then 3.5

OpenStudy (anonymous):

I don't have that on mine :(

OpenStudy (whpalmer4):

use google, like I suggested.

OpenStudy (anonymous):

lol I was just now.

OpenStudy (whpalmer4):

sorry, it's labeled \(x^y\) on the google keypad. just to the left of the 0

OpenStudy (anonymous):

1.22

OpenStudy (whpalmer4):

Yes! Multiply that by 3250 and you'll have the correct answer

OpenStudy (anonymous):

thanks for not giving up on me!

OpenStudy (anonymous):

$3965

OpenStudy (whpalmer4):

oh, I almost never do :-) it's frustrating how few students know how to use their calculators correctly, and unfortunately, the people most likely not to use their calculators correctly are also the ones most likely to not recognize that the answer they got doesn't make any sense. That's a bad combination :-)

OpenStudy (whpalmer4):

You should do it with a few more decimal places. Don't clear the result and type in 1.22 again, do the 1.06^3.5 and then just multiply by 3250

OpenStudy (whpalmer4):

correct answer is about $20 higher, and I'm sure Jose wants the extra $20 :-)

OpenStudy (anonymous):

My answer choices are $3932.50 $3985.23 $4752.00 $5200.00

OpenStudy (whpalmer4):

I see that google will let you just enter 3250*1.06(x^y)3.5 = and get the right result. back in my day, calculators didn't understand the proper order of execution, so you wouldn't necessarily get the right ansewr

OpenStudy (anonymous):

ok so just add it to the amount?

OpenStudy (whpalmer4):

no, do as I instructed. press AC to clear everything out, then enter \[3250\text{x}1.06x^y3.5=\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

$3985.23

OpenStudy (anonymous):

: /

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

isn't that one of your answer choices, down to the last penny?

OpenStudy (whpalmer4):

it is correct. you can check it by hand and eliminate 2 of the answers. after 1 year, 6% interest added to 3250 = 3250 + 0.06*3250 = 3445 after 2 years, another 6% interest added to 3445 = 3445 + 0.06*3445 = 3651.70 after 3 years, another 6% interest added to 3651.70 = 3651.70 + 0.06*3651.70 = 3870.80 now half a year's interest added to 3870.80 is going to be 3870.80 + 1/2*6%*3870.80 = 3986.92, which is very close to your result, certainly closer than any of the other answer choices